# Random Utility with Aggregation

### v2

@@ Appendix A Proofs
The necessity is trivial. We prove the sufficiency. Assume the two conditions and fix an aggregation function $X$ such that $|X(a_{0})|=|\mathscr{A}_{N}|+1$.
@@
By Partial-RU, there exists $\tilde{\mu}\in\Delta(\mathscr{L}(\mathscr{A}_{N}))$ that rationalizes $\rho$ on $2^{\mathscr{A}_{N}}\setminus\{\emptyset\}$. For each $y\in\mathscr{A}_{N}$, choose an element of $X(a_{0})$ denoted by $x_{0}(y)$. We will also fix a unobservable alternative $\underline{x}$ that plays a special role in the following. Note that such alternatives exist because $|X(a_{0})|=|\mathscr{A}_{N}|+1$.
@@
Let $\rho_{\scriptscriptstyle\mathscr{X}}$ be a random utility model that is rationalized by $\mu_{{\scriptscriptstyle\mathscr{X}}}$. Now fix $D\subseteq\mathscr{A}_{N}$. In the following, we will construct $\lambda\in\prod_{b\in D\cup a_{0}}(2^{X(b)}\setminus\{\emptyset\})$ to establish the following equality for all $b\in D\cup a_{0}$:
@@
Note that (13) is implied by (14) by the fact that the probability must sum to one. The purpose of our proof is, thus, to show the existence of $\lambda$ on $2^{X(a_{0})}\setminus\{\emptyset\}$ that satisfies the following equality for all $y\in D\subseteq\mathscr{A}_{N}$, [^18] Remember that this $\lambda$ is a composition distribution for the choice set $D\cup\{a_{0}\}$. For another choice set $E\cup\{a_{0}\}$, one can construct a desirable composition distribution in the same way.
@@
where the fourth equality holds because $\rho_{\scriptscriptstyle\mathscr{X}}(D,y_{0})=\rho(D,y_{0})$ because all elements in $D$ are non-aggregated categories.
@@
Note first that $\lambda_{n}$ is a probability measure given that $\lambda_{n-1}$ is a probability measure. Note also this modification of $\lambda_{n}$ from $\lambda_{n-1}$ does not change the chance that $y_{i}$ is chosen for any $i\leq n-1$; this only changes the chance that $y_{n}$ is chosen because adding $x_{0}(y_{n})$ changes only the choice frequency of $y_{n}$ as in (A.1). Moreover, we obtain the desired equalities for the choice frequency of $y_{n}$ as follows:
@@
$=c_{n}\lambda_{0}(\underline{x})\rho(D,y_{n})\quad(\because\sum_{S\in supp(\lambda_{n-1})\setminus\{X(a_{0})\}}\lambda_{n-1}(S)=1-\lambda_{n-1}(X(a_{0}))=1-\lambda_{0}(X(a_{0}))=\lambda_{0}(\{\underline{x}\}))$
@@
A.2 Proof of Remark 3.5
Six $\succ\in\mathscr{L}(\mathscr{A})$ and $\mathscr{E}\subseteq\mathscr{D}$. Choose any two elements of $X(a_{0})$ denoted by $\overline{x}$ and $\underline{x}$. Consider the following ranking $\succ^{\prime}\in\mathscr{L}(\mathscr{X})$ such that (i) $\succ$ and $\succ^{\prime}$ coincide on $\mathscr{A}_{N}$ (i.e., $a\succ b$ if and only if $x\succ^{\prime}y$ for all $a,b\in\mathscr{A}_{N}$, where $X(a)=\{x\}$ and $X(b)=\{y\}$); (ii) for any $y\in\mathscr{A}_{N}$, $\overline{x}\succ^{\prime}y\succ^{\prime}\underline{x}$. For all $D\in\mathscr{D}$, define $\lambda_{D}(\{\overline{x}\})=1$ for any $D\not\in\mathscr{E}$; and $\lambda_{D}(\{\underline{x}\})=1$ for any $D\in\mathscr{E}$. It is easy to see that $(\delta_{\succ^{\prime}},\lambda)$ RU- rationalizes $\rho^{\succ}_{\mathscr{E}}$. Note that this construction only relies on $|X(a_{0})|\geq 2$ which is implied by the outside option approach, thus Remark 3.5 does not depend on $X(a_{0})$.
A.3 Proof of Theorem 3.12
The necessity is trivial. We show the sufficiency. Fix an aggregation function $X$ such that $|X(a)|=|\mathscr{A}_{N}|+2$ for all $a\in\mathscr{A}_{A}$. For each $a\in\mathscr{A}_{A}$ and $y\in\mathscr{A}_{N}$, choose an element of $X(a)$, denoted by $x_{a}(y)$. We also fix two alternatives $\overline{x}_{a},\underline{x}_{a}\in X(a)$. Note that there exist such alternatives under the assumption that $|X(a)|=|\mathscr{A}_{N}|+2$ for all $a\in\mathscr{A}_{A}$. We again abuse notation and assume that each category $a\in\mathscr{A}_{N}$ is composed of itself, that is $X(a)=\{a\}$.
By Partial RU- rationality, there exists a distribution $\tilde{\mu}$ over linear orders $\tilde{\succ}$ on $\mathscr{A}_{N}$ that rationalizes $\rho$ on subsets of $\mathscr{A}_{N}$. Based on this distribution, we will construct a RUM $\rho$ on the entire set $\mathscr{X}$ of the underlying alternatives as follows. For each ranking on $\mathscr{A}_{N}$ in the support of $\tilde{\mu}$, we construct ranking on $\mathscr{X}$: we place $x_{a}(y)$ right above $y$, $\overline{x}_{a}$ and $\underline{x}_{a}$ at the bottom with $\overline{x}_{a}$s above $\underline{x}_{a}$s. The order among aggregated categories is arbitrarily fixed. For example, let $\{a_{0},a_{1}\}=\mathscr{A}_{A}$ and fix an order by index. Also let $\mathscr{A}_{N}=\{y_{0},y_{1}\}$ and $y_{0}\ \tilde{\succ}\ y_{1}$. Then, we define $\succ$ on $\mathscr{X}$ as follows:
Fix $D\subseteq\mathscr{A}_{N}$ and $E\subseteq\mathscr{A}_{A}$. In the following, we will construct $\lambda\in\prod_{b\in D\cup E}(2^{X(b)}\setminus\{\emptyset\})$ to establish the following equality for all $b\in D\cup E$:
With using the same simplified notation and the reasoning in the proof of Theorem 3.1, the purpose of the proof is to show the existence of a probability distribution $\lambda$ on $\prod_{b\in E}(2^{X(b)}\setminus\{\emptyset\})$ that satisfies
and
The proof consists of two steps. At the first step, for each $a\in E$ we will construct a $\lambda^{a}$ in the same way as the single outside option case (i.e., Theorem 3.1) except replace $\underline{x}$ with $\overline{x}_{a}$ and add $\underline{x}_{a}$ to all the sets. Given $\lambda^{a}$, in the second step, by taking a convex combination of $\lambda^{a}$, we will obtain the desired $\lambda$ that satisfy (16) and (17) for the given $D$ and $E$.
$=\rho(D\cup E,y_{0}).$
$=\rho(D\cup E,y_{n}),$
where the second equality holds because $x_{a}(y_{n})\succ y_{n}$ for all $\succ\in supp(\mu_{{\scriptscriptstyle\mathscr{X}}})$; the second to the last equality holds because $\sum_{(S_{b})_{b\in E}\in supp(\lambda^{a}_{n-1})\setminus(X(b))_{b\in E}}\lambda^{a}_{n-1}((S_{b})_{b\in E})=1-\lambda^{a}_{n-1}((X(b))_{b\in E})=1-\lambda^{a}_{0}((X(b))_{b\in E})=\lambda^{a}_{0}(\overline{x}_{a},(\underline{x})_{E\setminus a})$; and finally the third to the last equality holds because $y_{n}$ is a maximum element among $D$ with respect to $\succ$ if and only if $y_{n}$ is a maximum element among $D\cup(\cup_{b\in E}S_{b})$ with respect to $\succ$ (Note hat this equivalence holds because $S_{b}\subseteq\{x_{b}(y_{i})|i<n\}\cup\{\underline{x}_{b},\overline{x}_{b}|b\in\mathscr{A}_{A}\}$ and $x_{b}(y_{i})$ is the immediate predecessor of $y_{i}$ for all $\succ\in supp(\mu_{{\scriptscriptstyle\mathscr{X}}})$). It follows that $\lambda^{a}$ satisfies (16) as desired: for all $y\in D$
Step 2: In the following, given $(\lambda^{a})_{a\in\mathscr{A}_{A}}$, we will construct $\lambda$ that satisfies (17) as well as (16). First observe that for all $a\in\mathscr{A}_{A}$,
Since $\sum_{a\in E}\rho(D\cup E,a)=1-\sum_{y\in D}\rho(D\cup E,y)$ and each $\lambda^{a}$ is a probability distribution on $\prod_{a\in E}2^{X(a)}\setminus\{\emptyset\}$, so is $\lambda$. Moreover, we have the desired equalities for choice frequency of each aggregated category $b\in E$ as follows:

### v3

@@ Appendix D Additional theoretical results
The necessity is trivial. We prove the sufficiency. Assume the two conditions and fix an aggregation correspondence $X$ such that $|X(a_{0})|=|\mathscr{A}_{A}|+1$.
@@
By Partial-RU, there exists $\tilde{\mu}\in\Delta(\mathscr{L}(\mathscr{A}_{A}))$ that rationalizes $\rho$ on $2^{\mathscr{A}_{A}}\setminus\{\emptyset\}$. For each $y\in\mathscr{A}_{A}$, choose an element of $X(a_{0})$ denoted by $x_{0}(y)$. We will also fix a unobservable alternative $\underline{x}$ that plays a special role in the following. Note that such alternatives exist because $|X(a_{0})|=|\mathscr{A}_{A}|+1$.
@@
Let $\rho_{\scriptscriptstyle\mathscr{X}}$ be a random utility model that is rationalized by $\mu_{{\scriptscriptstyle\mathscr{X}}}$. Now fix $D\subseteq\mathscr{A}_{A}$. In the following, we will construct $\lambda\in\prod_{b\in D\cup a_{0}}(2^{X(b)}\setminus\{\emptyset\})$ to establish the following equality for all $b\in D\cup a_{0}$:
@@
Note that (18) is implied by (19) by the fact that the probability must sum to one. The purpose of our proof is, thus, to show the existence of $\lambda$ on $2^{X(a_{0})}\setminus\{\emptyset\}$ that satisfies the following equality for all $y\in D\subseteq\mathscr{A}_{A}$, [^33] Remember that this $\lambda$ is a composition distribution for the choice set $D\cup\{a_{0}\}$. For another choice set $E\cup\{a_{0}\}$, one can construct a desirable composition distribution in the same way.
@@
where the fourth equality holds because $\rho_{\scriptscriptstyle\mathscr{X}}(D,y_{0})=\rho(D,y_{0})$ because all elements in $D$ are atomic aggregates.
@@
Note first that $\lambda_{n}$ is a probability measure given that $\lambda_{n-1}$ is a probability measure. Note also this modification of $\lambda_{n}$ from $\lambda_{n-1}$ does not change the chance that $y_{i}$ is chosen for any $i\leq n-1$; this only changes the chance that $y_{n}$ is chosen because adding $x_{0}(y_{n})$ changes only the choice frequency of $y_{n}$ as in (D.1). Moreover, we obtain the desired equalities for the choice frequency of $y_{n}$ as follows:
@@
$=c_{n}\lambda_{0}(\underline{x})\rho(D,y_{n})$
$(\because\sum_{S\in supp(\lambda_{n-1})\setminus\{X(a_{0})\}}\lambda_{n-1}(S)=1-\lambda_{n-1}(X(a_{0}))=1-\lambda_{0}(X(a_{0}))=\lambda_{0}(\{\underline{x}\}))$
@@
D.2 Technical remarks
Remark D.1.
For Theorem 3.1, in order to make sense of Limited Monotonicity, we need the following assumption of the domain of $\rho$: if $\rho$ is defined on $(D\cup\{a_{0}\},\cdot)$, then $\rho$ is defined on $(D,\cdot)$ (i.e., if $D\cup\{a_{0}\}\in\mathscr{D}$, then $D\in\mathscr{D}$). In this case, Limited Monotonicity should be defined as: for all $D\subseteq\mathscr{A}_{A}$ such that $D\cup\{a_{0}\}\in\mathscr{D}$, $\rho$ satisfies $\rho(D,a)\geq\rho(D\cup\{a_{0}\},a)$ for all $a\in D$.
Remark D.2.
In our notion of RU rationality, we assume that the composition of aggregates and the agent’s preferences are independent. A natural extension is to allow the composition distribution to depend on the agent’s preferences. That is, a composition distribution $\lambda_{\succ,A}$ is a distribution over  $\prod_{a\in A}(2^{X(a)}\setminus\{\emptyset\})$ that may jointly depend on the available aggregates $A$ and the agent’s preference $\succ\in\mathscr{L}(\mathscr{X})$. For example, we may allow the case where consumers with a preference for high quality goods tend to have a higher quality outside option. [^34] We do not, however, allow the composition of the outside option to enter into the consumer’s underlying choices. This type of dependence would lead to RUM violations in the underlying choice which would lead to a model with no observable implications.
We say that $\rho$ is General RU rational if there exists $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda_{\succ,A})$ such that $\rho(A,a)$ equals
$\sum_{(S_{b})_{b\in A}\in\underset{b\in A}{\prod}(2^{X(b)}\setminus\{\emptyset\})}\sum_{\succ\in\mathscr{L}(\mathscr{X})}1(\exists x\in S_{a},\ \forall y\in\underset{b\in A\setminus\{a\}}{\cup}S_{b},\ x\succ y))\lambda_{\succ,A}((S_{b})_{b\in A})\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ).$ (21)
Remark D.3.
That the threshold equals $|\mathscr{A}_{A}|+1$ can be seen from how weak Limited Monotonicity is: it only implies $\rho(D,b)\ \geq\ \rho(D\cup\{a_{0}\},b)\quad\text{for all }b\in\mathscr{A}_{A}$, and therefore allows essentially arbitrary (and highly heterogeneous) substitution across atomic aggregates. For instance, it permits $\rho(D,a)=\rho(D\cup\{a_{0}\},a)$ for some $a$ (no substitution) while $\rho(D,b)\gg\rho(D\cup\{a_{0}\},b)$ for other $b$ (strong substitution). To rationalize such patterns with an underlying RUM, one needs, for each $b\in\mathscr{A}_{A}$, a distinct underlying alternative that can become available via $a_{0}$ and is ranked just above $b$, yielding $|\mathscr{A}_{A}|$ underlying alternatives. If $\rho(D,a)=\rho(D\cup\{a_{0}\},a)$ holds for some $a$, then $a_{0}$ must be worse than every atomic aggregate, which requires one additional underlying alternative ranked below all of $\mathscr{A}_{A}$. Hence the threshold is $|\mathscr{A}_{A}|+1$.
Remark D.4.
We assumed the full domain $\mathscr{D}=2^{\mathscr{A}}\setminus\{\emptyset\}$ for Theorem 3.3. However, the proof only requires that there exists a labeling $\mathscr{A}_{A}=\{y_{1},...,y_{|\mathscr{A}|}\}$ such that for all $1\leq n\leq|\mathscr{A}|$, the sets $D_{n}=\{y_{1},...,y_{n}\}$ and $D_{n}\cup\{a_{0}\}$, as well as any other $D\subset\mathscr{A}_{A}$ are in $\mathscr{D}$,.
Remark D.5.